The total electric flux through a closed cylindrical (length = 1.2 m, diameter = 0.20 m) surface is
equal to 5.0 N  m2
/C. Determine the net charge within the cylinder
1
Expert's answer
2020-11-17T07:13:10-0500

From the Gauss’ law the net electric flux of a closed object is

"Φ = \frac{q}{ε_0}"

q is the charge within the cylinder

ε0 is permittivity in free space

"q = Φ \times ε_0"

"q = (-5.0\; N\;m^2/C)(8.55 \times 10^{-12}\;C^2/Nm^2)"

"q = -44.25 \times 10^{-12}\;C = -44 \;pC"